OpenStudy (htmerritt):

Three teammates had different point totals at the girls' basketball game. To determine the number of points Effie had, multiply Toni's points by 3, subtract 8, and then multiply the difference by 2. To determine the number of points Linda had, add 9 to Toni's points and divide the sum by 3. How many points did each girl have if Eiffe scored 9 more than Toni and Linda combined?

OpenStudy (aum):

Let T be Toni's points; E be Effie's points; L be Linda's points. "To determine the number of points Effie had, multiply Toni's points by 3, subtract 8, and then multiply the difference by 2.": E = (3*T - 8) * 2 = 6T - 16 E = 6T - 16 ---- (1) "To determine the number of points Linda had, add 9 to Toni's points and divide the sum by 3.": L = (T + 9) / 3 ----- (2) "Eiffe scored 9 more than Toni and Linda combined": E = T + L + 9 ----- (3) Three equations and three unknowns. Solve them.

OpenStudy (aum):

Substitute E from (1) and L from (2) into (3). Solve for T. Then sub T in (2) and solve for L. Then sub T and L into (3) and solve for E.